OpenStudy (anonymous):

Show the conjecture is false by finding a counterexample. 1. the quotient of the two whole numbers is a whole number 2. the difference of the abosolute value of the two numbers is positive, meaning... |a| - |b| >0 3. if m≠ -1 then m *over* m+1 <1

OpenStudy (anonymous):

do you know what "counter example" means?

OpenStudy (anonymous):

1) 3/8 = 0.375 2) abs(1) - abs(2) = 1 - 2 = -1 < 0

OpenStudy (anonymous):

actually no because the teacher kinda just gives us the homework its her first year teaching and she has no clue what shes doing... and were all extremely behind..

OpenStudy (anonymous):

ok so we go slow

OpenStudy (anonymous):

counter example: say that what you are trying to prove false is true, then use that to find an example to prove it false. (bascially)

OpenStudy (anonymous):

the quotient of the two whole numbers is a whole number a counter example would be two whole numbers whose quotient is NOT a whole number

OpenStudy (anonymous):

amost any two whole numbers will do not \(12\) and \(4\) because \(\frac{12}{4}=3\) IS a whole number, but (12\) and \((5\) would work because \[\frac{12}{5}=2.4\]is NOT a whole number

OpenStudy (anonymous):

for the difference of the abosolute value of the two numbers is positive, meaning... |a| - |b| >0 you have to come up with two numbers \(a,b\) where \(|a|-|b|\) is NOT \(>0\) in other words, two numbers \(a,b\) where \(|a|-|b|<0\)

OpenStudy (anonymous):

that should not be very hard, just pick two numbers where \(b\) is bigger than \(a\)

OpenStudy (anonymous):

Okay thank you so much... I'd be lost without openstudy because she expects us to know it after she goes over it once and then when we ask she gets mad at us.

OpenStudy (anonymous):

you good with the last one too?

OpenStudy (anonymous):

I think so...

OpenStudy (anonymous):

what do you need to find? you need to find some number \(m\) so that \[\frac{m}{m+1}>1\] this might be tricky, maybe not

OpenStudy (anonymous):

whole numbers will not work

OpenStudy (anonymous):

So find a number thats smaller than it?

OpenStudy (anonymous):

find some number \(m\) where \(\frac{m}{m+1}>1\) \(5\) will not work, because \[\frac{5}{5+1}=\frac{5}{6}<1\]

OpenStudy (anonymous):

so you're saying it will come out to be a decimal but in fraction form?

OpenStudy (anonymous):

no let me say it in english you need to pick a number, so what when you divide that number by a number that is bigger by 1, you get a result that is LARGER than one, not smaller would you like me to give you an example?

OpenStudy (anonymous):

or a hint?

OpenStudy (anonymous):

yes please...

OpenStudy (anonymous):

pick a number that is negative any number less than \(-1\) and we can try it

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

ok so \(m=-2\) and then \(m+1=-2+1=-1\) right?

OpenStudy (anonymous):

then \[\frac{m}{m+1}=\frac{-2}{-2+1}=\frac{-2}{-1}=2\]

OpenStudy (anonymous):

and there you have it, a COUNTEREXAMPLE, because \(2>1\) not less than

OpenStudy (anonymous):

okay thank you so much

OpenStudy (anonymous):

yw that all?

OpenStudy (anonymous):

yes you've helped so much thank you

OpenStudy (anonymous):

your welcome